OpenStudy (anonymous):

A ball is thrown from an initial height of 1 meter with an initial upward velocity of 25/ms.The ball's height h (in meters) after t seconds is given by the following. h=1+25t-5t^2 Find all values of t for which the ball's height is 6 meters. Round your answer(s) to the nearest hundredth.

OpenStudy (koikkara):

@kamelish Can you draw the picture and write down the given data, like \(h=?\) etc ...

OpenStudy (anonymous):

it wont let me @Koikkara

OpenStudy (phi):

\[ h=1+25t-52t^2 \] ball's height is 6 . That means h is 6. \[ 6 = 1+25t-52t^2 \] put that in standard form \[ -52t^2 +25t -5=0 \] It looks like we should use the quadratic formula to solve for t do you know the quadratic formula?

OpenStudy (anonymous):

its 5t^2 i made a mistake @phi

OpenStudy (phi):

that's good, a much nicer number \[ -5t^2 +25t+1=6 \\ -5t^2 + 25t -5= 0 \] divide both sides by -5 (to simplify the equation a little bit) \[ t^2 -5t+1=0\] but you still need to use the quadratic formula. Do you know it ?

OpenStudy (anonymous):

i know it a little bit @phi

OpenStudy (anonymous):

t=4.79 or t=0.20? @phi

OpenStudy (phi):

if you have multiple choice, you can test each answer in the original formula \[ 1+25t-5t^2 = 6 \] and see if you get 6 for the left-hand side

OpenStudy (anonymous):

thnaks =) @phi